# Plenoptic Cameras; Light Field; Calibration; Refocusing; Simulation

### 20 Feb 2024 v1

@@ 1 Introduction
Plenoptic cameras as initially described by Lippmann Lippmann (1908) and Ives Ives (1928) combine a traditional camera with an additional microlens array (MLA) located between the main lens and the sensor. Over the years, two primary designs have been extensively studied and brought to market, the standard plenoptic camera (SPC) Adelson and Wang (1992) Ng et al. (2005) and the focused plenoptic camera (FPC) Georgiev and Intwala (2006)Perwass and Wietzke (2012), which mainly differ in the microlens focus distance. Due to the earlier commercialization, a larger angular resolution and a simpler decoding process, the SPC still remains popular, even though it has certain disadvantages in terms of the spatial resolution and depth of field when compared to the multi-focus variant of the FPC Perwass and Wietzke (2012). Classical applications for SPCs include depth reconstruction Adelson and Wang (1992) and post-capture refocusing from single shots Ng et al. (2005) and as a first step to achieve these, the raw 2D image of a plenoptic camera is usually de-multiplexed and resampled into a 4D light field Dansereau et al. (2013) as shown in figure 1. For this re-parametrization procedure the knowledge about the exact position of each microlens image center (MIC) is crucial, as any inaccuracies in their locations can result in computational errors affecting the quality of the refocused images Hahne et al. (2018). Futhermore, a formal connection between the MICs and the plenoptic camera optical setup is required in order to relate the light field within the camera to the optical reality outside the camera, e.g. for finding the correct refocusing parameters for the desired object distance Dansereau et al. (2013)Hahne et al. (2018).
@@
In this context, our work aims to highlight the importance of the exit pupil again. To this end first a paraxial model of the SPC under consideration of the exit pupil is described, directly relating the refocusing shift Ng et al. (2005) to the object distance. The expected errors of the models ignoring the exit pupil is formally analyzed and later verified through a ray-tracing-based simulation of various plenoptic cameras using real lens data. Subsequently, multiple works in the domain of plenoptic camera calibration are revisited and examined with respect to the need for a more complex lens model. More specifically, first the popular work of Dansereau et al. Dansereau et al. (2013) is revisited as well as the works of Zhang et al. Zhang et al. (2018) and Monteiro et al. Monteiro et al. (2019) building upon Dansereau’s ideas. In these cases, one can conclude that the parameters of the respective calibration models are sufficiently general to permit the simplicity of a main lens model without considering the exit pupil. However, this does only hold true since these works do not require a specific interpretation of the model parameters. For the work of Pertuz et al. Pertuz et al. (2018) on the other hand, which also employs the decoding from Dansereau et al. (2013) for metric distance measurement, it is shown, that the oversimplified main lens model leads to an incorrect interpretation of their metric refocusing model parameters. In summary, our contributions are:
@@
  • Publicly available [^1] https://gitlab.com/ungetym/blender-camera-generator SPC designs and a new ray-tracing framework supporting a large data base of lens designs and enabling a quick generation of new plenoptic camera setups.
@@ 1.1 Related Work
Later the focused plenoptic camera (FPC) was presented by Lumsdaine and Georgiev Lumsdaine and Georgiev (2009) and extended by Perwass and Wietzke Perwass and Wietzke (2012) to feature multi-focal MLAs. For the FPC instead of directly dissecting the scene into its directional components the microlenses are focused at the scene’s virtual image inside the camera. This arrangement is advantageous in maintaining a greater portion of the conventional camera’s spatial resolution albeit at the expense of a decrease in angular resolution compared to the SPC.
Our work mainly deals with SPCs since the problems arising from an ignored exit pupil are more directly observable in this case. Note, however, that this work is intended to be the base of further research also improving FPC calibration algorithms suffering from the same model flaws.
@@
SPC calibration: The calibration of plenoptic cameras plays a crucial role in relating the captured light field within the camera to the 3D world in front of the camera. To this end, Dansereau et al. Dansereau et al. (2013) present a method for demultiplexing the raw image of an SPC into a 2D array of sub-aperture images and using these for the geometric calibration. Due to its popularity and accessibility in form of a Matlab toolbox, this work is still being used as base for publications concerning the processing of SPC data. With respect to calibration models Zhang et al. Zhang et al. (2018) and Monteiro et al. Monteiro et al. (2019) modify the ideas of Dansereau et al. (2013) in order to associate a plenoptic camera with an equivalent multi-camera array. Both works make direct use of the decoding process proposed by Dansereau et al. Dansereau et al. (2013). Pertuz et al. Pertuz et al. (2018) also follow this approach and propose a focus-based metric depth estimation. A different calibration approach, directly using line-features in the microlens images, is presented by Bok et al. Bok et al. (2016). And Hahne et al. Hahne et al. (2016)Hahne et al. (2018) describe the refocusing distance based on known main lens and MLA parameters under consideration of the exit pupil. Schambach et al. Schambach and León (2020) further improve the MIC detection and Mignard-Debise and Ihrke Mignard-Debise and Ihrke (2019) analyze the effect of vignetting on calibration models.
@@
While especially Hahne et al. Hahne et al. (2016) consider the exit pupil and its connection to the microlens image geometry in a similar fashion to this work, this is neither put into the direct context of pre-existing works by a comparative evaluation nor does it provide an analysis of the expected errors resulting from oversimplified lens models.
@@
Lens models and simulation: In the domain of ray tracing-based camera simulation, realistic main lens models which consider all lens components and their respective properties have been used for over two decades, either explicitely by direct modeling as in Kolb et al. Kolb et al. (1995) and Wu et al. Wu et al. (2011) or implicitly via learned black-box lens models as proposed in Zheng et al. Zheng and Zheng (2017). Regarding plenoptic cameras, most previous work uses oversimplified models for rendering, such as pinhole cameras or multi camera arrays modeling the MLA, but without a model for the main lens Fleischmann and Koch (2014)Zhang et al. (2015)Liang and Ramamoorthi (2015). More recently, Nürnberg et al. Nürnberg et al. (2019) as well as our group Michels et al. (2018) provided simulations of plenoptic cameras without oversimplifying the main lens. Due to familiarity we extended our previous work for the synthetic experiments.
@@ 2.2 Preliminaries - Light Field Parametrization
Despite being a standard tool when working with light field data, the two-plane parametrization as described by Levoy and Hanrahan Levoy and Hanrahan (2023) and used in various popular works including the work of Dansereau et al. Dansereau et al. (2013) and Ng et al. Ng et al. (2005) are reiterated in this section for two reasons. First, the previous descriptions do not consider the exit pupil, and second, the literature is not consistent in terms of the underlying data representation. While Dansereau et al. (2013) uses the raw camera image, indexed by integer pixel coordinates, to base the description on, Ng et al. (2005) assumes known metric coordinates for every pixel. We follow the former approach to facilitate the reproduction of our results.
Given an SPC following the thick lens model with an exit pupil as visualized in figure 2, the light field inside this camera can be parameterized using two planes - the MLA and the exit pupil plane where the former serves as virtual sensor plane and the latter can be interpreted as virtual lens plane. By following the decoding process of Dansereau et al. Dansereau et al. (2013) the 4D light field can be parametrized as $L_{F}(i,j,k,l)$ with integer indices $(k,l)$ for the uniformly sampled sub-aperture image and $(i,j)$ for the pixel coordinates in that image.
The corresponding metric parametrization $\tilde{L}_{F}(s,t,u,v)$ describes the intensity of light captured at the MLA plane point $(s,t,d)$ coming from the exit plane point $(u,v,X)$. In accordance with Ng et al. Ng et al. (2005)Ng (2005) 222 Note, that Ng et al. implicitly assume $X=0$ since the main lens in that work is modeled by a thin lens. While this is approximately correct for the two tested main lenses in Ng et al. (2005), a Zeiss Planar T* 2/80 with $X=7.8mm=0.098\cdot f_{M}$ and a Zeiss Sonnar T* 2.8/140 with $X=-4mm=-0.029\cdot f_{M}$ (compare zei (a)zei (b)), section 3 shows, that this assumption does not hold in general. the distance between these two parametrization planes is denoted as $F:=d-X$. To calculate the metric parametrization from a given integer parametrization, note, that the pixel pitch $\Delta_{ST}$ of the virtual sensor, i.e. the step size in the $ST$-Plane, corresponds to the microlens pitch, i.e. $\Delta_{ST}=d_{\text{ML}}$ and, as shown in figure 3, the step size in the virtual lens plane, i.e. the $UV$-plane, can be calculated by means of the triangle equality as $\Delta_{UV}=\frac{s_{px}\cdot F}{f_{m}}$. With these step sizes the light field parametrized in metric coordinates $(s,t,u,v)$ is given by
@@ 3 Error Analysis
$ERR_{S}(\textbf{o}):=\frac{S(\textbf{o})-\tilde{S}(\textbf{o})}{S(\textbf{o})}=X\cdot\frac{\textbf{o}(d-f_{M})-d\cdot f_{M}}{\textbf{o}\cdot f_{M}(d-X)}$ (11)
@@
$ERR_{S}(\lambda):=ERR_{S}(\lambda\cdot\textbf{o}_{f})=\frac{X(\lambda-1)}{\lambda\textbf{o}_{f}\cdot\left(1-\frac{X}{d}\right)}.$ (12)
@@
$ERR_{\tilde{\textbf{o}}}(\textbf{o}):=\frac{\textbf{o}-\tilde{\textbf{o}}(S(\textbf{o}))}{\textbf{o}}=\frac{X(d-f_{M})+\frac{\textbf{o}}{\textbf{o}_{f}}X(f_{M}-d)+Xf_{M}d\left(\frac{1}{\textbf{o}_{f}}-\frac{1}{\textbf{o}}\right)}{f_{M}(d-X)+\frac{\textbf{o}}{\textbf{o}_{f}}X(f_{M}-d)+Xf_{M}d\frac{1}{\textbf{o}_{f}}}$ (13)
@@
$ERR_{\tilde{\textbf{o}}}(\lambda):=ERR_{\tilde{\textbf{o}}}(\lambda\textbf{o}_{f})=\frac{X(\lambda-1)^{2}}{X(\lambda-1)\lambda-\lambda\textbf{o}_{f}\left(1-\frac{X}{d}\right)}.$ (14)
@@
$ERR_{\tilde{S}}(\textbf{o}):=\frac{\textbf{o}-\textbf{o}(\tilde{S}(\textbf{o}))}{\textbf{o}}=1-\frac{f_{M}\mathbf{\Delta}-f_{M}X\tilde{\mathbf{\Delta}}\left(\frac{1}{\textbf{o}}-\frac{1}{\textbf{o}_{f}}\right)}{\tilde{\mathbf{\Delta}}\left(1-\frac{\textbf{o}}{\textbf{o}_{f}}\right)(f_{M}-X)+\mathbf{\Delta}f_{M}\frac{\textbf{o}}{\textbf{o}_{f}}}$ (15)
@@
$ERR_{\tilde{S}}(\lambda):=ERR_{\tilde{S}}(\lambda\textbf{o}_{f})=\frac{X(\lambda-1)^{2}}{\lambda^{2}X+\lambda\textbf{o}_{f}\left(1-\frac{X}{f_{M}}\right)}.$ (16)
@@
Figure 5: Left: Relative shift error. Mid/right: The two cases of relative object distance errors for the assumed camera focused at a finite distance which is met for a relative distance of $\frac{\textbf{o}}{\textbf{o}_{f}}=1$.
The visualization shows that all errors diverge for $\lambda\to 0$ with a rate depending on the positional relationship between the exit pupil and the principal plane $H_{\text{cam}}$. Beyond the focus distance at $\lambda=1$ the object errors again diverge while the shift error converges according to $ERR_{S}(\lambda)\underset{\lambda\to\infty}{\longrightarrow}\frac{X}{\textbf{o}_{f}(1-X/d)}$. Note, that these graphs present an ideal refocusing case free of aliasing artifacts and limiting optical properties such as the depth of field. Hence, later experiments only verify a section of these results within the respective physical and image processing limits.
In summary these examples show a large deviation between the esimated refocusing distances in models with and without consideration of the exit pupil whenever there is a non-zero distance $X$ between the exit pupil and $H_{\text{cam}}$. This leads to the question, how prevalent a significant $X\neq 0$ is in off-the-shelf main lenses. To answer this question, the data of 866 DSLR lenses listed by Claff Claff was collected and $X$ as well as $f_{M}$ were calculated via paraxial ray-tracing for each lens.

### 05 Apr 2024 v2

@@ 1 Introduction
Plenoptic cameras as initially described by Lippmann Lippmann (1908) and Ives Ives (1928) combine a traditional camera with an additional microlens array (MLA) located between the main lens and the sensor. Over the years, two primary designs have been extensively studied and brought to market, the standard plenoptic camera (SPC) Adelson and Wang (1992) Ng et al. (2005) and the focused plenoptic camera (FPC) Georgiev and Intwala (2006)Perwass and Wietzke (2012), which mainly differ in the microlens focus distance. Due to the earlier commercialization, a larger angular resolution and a simpler decoding process, the SPC still remains popular, even though it has certain disadvantages in terms of the spatial resolution and depth of field when compared to the multi-focus variant of the FPC Perwass and Wietzke (2012). Classical applications for SPCs include depth reconstruction Adelson and Wang (1992) and post-capture refocusing from single shots Ng et al. (2005) and as a first step to achieve these, the raw 2D image of a plenoptic camera is usually de-multiplexed and resampled into a 4D light field Dansereau et al. (2013) as shown in figure 1. For this reparametrization procedure the knowledge about the exact position of each microlens image center (MIC) is crucial, as any inaccuracies in their locations can result in computational errors affecting the quality of the refocused images Hahne et al. (2018). Furthermore, a formal connection between the MICs and the plenoptic camera optical setup is required in order to relate the light field within the camera to the optical reality outside the camera, e.g. for finding the correct refocusing parameters for the desired object distance Dansereau et al. (2013)Hahne et al. (2018).
@@
In this context, our work aims to highlight the importance of the exit pupil again. To this end first a paraxial model of the SPC under consideration of the exit pupil is described, directly relating the refocusing shift Ng et al. (2005) to the object distance. The expected errors of the models ignoring the exit pupil is formally analyzed and later verified through a ray-tracing-based simulation of various plenoptic cameras in Blender ble using real lens data. Subsequently, multiple works in the domain of plenoptic camera calibration are revisited and examined with respect to the need for a more complex lens model. More specifically, first the popular work of Dansereau et al. Dansereau et al. (2013) is revisited as well as the works of Zhang et al. Zhang et al. (2018) and Monteiro et al. Monteiro et al. (2019) building upon Dansereau’s ideas. In these cases, one can conclude that the parameters of the respective calibration models are sufficiently general to permit the simplicity of a main lens model without considering the exit pupil. However, this does only hold true since these works do not require a specific interpretation of the model parameters. For the work of Pertuz et al. Pertuz et al. (2018) on the other hand, which also employs the decoding from Dansereau et al. (2013) for metric distance measurement, it is shown, that the oversimplified main lens model leads to an incorrect interpretation of their metric refocusing model parameters. In summary, our contributions are:
@@
  • Publicly available [^1] https://gitlab.com/ungetym/SPC-revisited and https://gitlab.com/ungetym/blender-camera-generator SPC designs and a camera simulation framework based on Blender ble supporting a large data base of lens designs and enabling a quick generation of new plenoptic camera setups.
@@ 1.1 Related Work
Later, the focused plenoptic camera (FPC) was presented by Lumsdaine and Georgiev Lumsdaine and Georgiev (2009) and extended by Perwass and Wietzke Perwass and Wietzke (2012) to feature multifocal MLAs for an extended depth of field. For the FPC, instead of directly dissecting the scene into its directional components, the microlenses are focused at the scene’s virtual image inside the camera. This arrangement is advantageous in maintaining a greater portion of the conventional camera’s spatial resolution albeit at the expense of a decrease in angular resolution compared to the SPC.
Despite the advantages of FPCs especially when combined with multifocal MLAs, we chose to focus this work on the SPC due to the straightforward image processing pipeline. More specific, to show the effect of the exit pupil, the application of post-shot refocusing is used throughout this work. And while this is a possible application for an FPC, the process involves a direct Wanner and Goldluecke (2012) or indirect Georgiev and Lumsdaine (2010) depth estimation which can introduce artifacts affecting subsequent processing steps. In contrast, for a well configured SPC the post-shot refocusing only comprises of a demultiplexing step Dansereau et al. (2013) for the raw image with a subsequent shift-and-sum procedure for the resulting sub-aperture images. The simplicity of this pipeline reduces the quantity of artifacts resulting from complex interpolation and optimization steps and therefore allows a more direct analysis of the exit pupil’s effects. Note, however, that this work is intended to be the starting point for further research which will also analyze and improve FPC calibration algorithms suffering from the same model flaws as their SPC pendants.
@@
SPC calibration: The calibration of plenoptic cameras plays a crucial role in relating the captured light field within the camera to the 3D world in front of the camera. To this end, Dansereau et al. Dansereau et al. (2013) present a method for demultiplexing the raw image of an SPC into a 2D array of sub-aperture images and using these for the geometric calibration. Due to its popularity and accessibility in form of a Matlab toolbox, this work is still being used as base for publications concerning the processing of SPC data. With respect to calibration models Zhang et al. Zhang et al. (2018) and Monteiro et al. Monteiro et al. (2019) modify the ideas of Dansereau et al. (2013) in order to associate a plenoptic camera with an equivalent multi-camera array. Both works make direct use of the decoding process proposed by Dansereau et al. Dansereau et al. (2013). Pertuz et al. Pertuz et al. (2018) also follow this approach and propose a focus-based metric depth estimation.
Due to the popularity of the demultiplexing process of Dansereau et al. (2013), the revision of previous literature in section 4 focuses on this method and the approaches based on it. Nevertheless, there is further work, which in part also reduces the main lens to a thin lens and thereby ignores the effects of the exit pupil. One such calibration approach, directly using line-features in the microlens images, is presented by Bok et al. Bok et al. (2016) and with a similar model Zhao et al. Zhao et al. (2020) perform an SPC calibration based on plenoptic disc features. Thomasen et al. Thomason et al. (2014) as well as Suliga and Wrona Suliga and Wrona (2018) directly estimate the MLA pose and microlens pitch, but do not relate the captured light field to the scene-side light field. Like the approaches related to Dansereau et al. (2013), these works also assume the microlens images centers to be projections of the main lens center, i.e. the camera-side principal plane’s center. On the other hand Hahne et al. Hahne et al. (2018)Hahne et al. (2016) describe the refocusing distance based on known main lens and MLA parameters under consideration of the exit pupil. And further improvements to aspects of the calibration pipeline, which also acknowledge the exit pupil, are presented by Schambach et al. Schambach and León (2020) increasing the MIC detection accuracy and Mignard-Debise and Ihrke Mignard-Debise and Ihrke (2019) analyzing the effect of vignetting on calibration models.
@@
Of these works especially Hahne et al. Hahne et al. (2016) consider the exit pupil and its connection to the microlens image geometry in a similar fashion to this work, but this is neither put into direct context of pre-existing calibration methods by a comparative evaluation nor does it provide an analysis of the expected errors resulting from oversimplified lens models. Nevertheless, to further validate our model, which establishes the refocusing distance in terms of the two-plane parametrization, its equivalence to the chief ray intersection model in Hahne et al. (2016) is formally proven in section 4.1.
FPC calibration: Despite this work focusing on SPCs as explained above, there is also relevant work in the related field of FPC calibration. Johannsen et al. Johannsen et al. (2013) describe a metric reprojection model for FPCs incorporating a radial distortion model. This is enhanced by Heinze et al. Heinze et al. (2016) to also include the tilt and shift of the main lens as well as multi-focus MLAs. Further improvements to the distortion model are presented by Zeller et al. Zeller et al. (2016). All these approaches are based on the reconstruction of the virtual scene between MLA and main lens and associate these virtual 3d points to the known scene points. In contrast, Noury et al. Noury et al. (2017) propose an approach directly working on the microlens images, i.e. associating the scene points with their projections on the sensor without the intermediate step of calculating virtual depths. This method, however, is limited to single focus FPCs and models the microlenses as simple pinholes. Nousias et al. Nousias et al. (2017) on the other hand feature a more complete microlens model and directly include the estimation of multiple microlens focal lengths into their approach. And Wang et al. Wang et al. (2018) present a two-step model including a forward projection from the scene into the camera and a second projection from the virtual image to the sensor. More recently, Labussiere et al. Labussière et al. (2022) proposed a simultaneous calibration of the different microlens types in a multi-focus plenoptic camera by incorporating defocus blur into the features used for the parameter optimization.
None of the listed methods for FPC calibration directly consider the exit pupil and while most of these works, which require the identification of MICs, incorporate a scaling between the grid of microlens centers and the grid of MICs, this is usually a result of projecting the main lens center, i.e. the center of the camera-side principal plane, through the microlens centers. However, as observed by Hahne et al. Hahne et al. (2018)Hahne et al. (2016) for SPCs and confirmed in section 5, the MICs result from a projection of the exit pupil’s center instead. Thus using the distance between the simplified main lens plane and the MLA for the image formation model as well as the calculation of MICs could inadvertently reduce the degrees of freedom of the model. And while this might be desired in terms of an increased stability during the parameter optimization, this reduction should also be analyzed for FPC models. However, for reasons of clarity and comprehensibility we decided against also including the topic of FPC calibration into this work and leave this for future work.
@@
Lens models and simulation: In the domain of ray tracing-based camera simulation, realistic main lens models which consider all lens components and their respective properties have been used for over two decades, either explicitly by direct modeling as in Kolb et al. Kolb et al. (1995) and Wu et al. Wu et al. (2011) or implicitly via learned black-box lens models as proposed in Zheng et al. Zheng and Zheng (2017). Regarding plenoptic cameras, most previous work uses oversimplified models for rendering, such as pinhole cameras or multi camera arrays modeling the MLA, but without a model for the main lens Fleischmann and Koch (2014)Zhang et al. (2015)Liang and Ramamoorthi (2015). More recently, Nürnberg et al. Nürnberg et al. (2019) as well as our group Michels et al. (2018) provided simulations of plenoptic cameras without oversimplifying the main lens. Due to familiarity we extended our previous work for the synthetic experiments.
@@ 2.2 Preliminaries - Light Field Parametrization
Despite being a standard tool when working with light field data, the two-plane parametrization as described by Levoy and Hanrahan Levoy and Hanrahan (2023) and used in various popular works including the work of Dansereau et al. Dansereau et al. (2013) and Ng et al. Ng et al. (2005) are reiterated in this section for two reasons. First, the previous descriptions do not consider the exit pupil, and second, the literature is not consistent in terms of the underlying data representation. While Dansereau et al. (2013) uses the raw camera image, indexed by integer pixel coordinates, to base the description on, Ng et al. (2005) assumes known metric coordinates for every pixel. We follow the approach of Dansereau et al. (2013) to facilitate the reproduction of our results.
Given an SPC following the thick lens model with an exit pupil as visualized in figure 2, the light field inside this camera can be parameterized using two planes - the MLA, which serves as virtual sensor plane, and the exit pupil plane, which can be interpreted as virtual lens plane. By following the decoding process of Dansereau et al. Dansereau et al. (2013) the 4D light field can be parametrized as $L_{F}(i,j,k,l)$ with integer indices $(k,l)$ for the uniformly sampled sub-aperture image and $(i,j)$ for the pixel coordinates in that image.
The corresponding metric parametrization $\tilde{L}_{F}(s,t,u,v)$ describes the intensity of light captured at the MLA plane point $(s,t,d)$ coming from the exit plane point $(u,v,X)$. In accordance with Ng et al. Ng et al. (2005)Ng (2005) 222 Note, that Ng et al. implicitly assume $X=0$ since the main lens in that work is modeled by a thin lens. While this is approximately correct for the two tested main lenses in Ng et al. (2005), a Zeiss Planar T* 2/80 with $X=7.8mm=0.098\cdot f_{M}$ and a Zeiss Sonnar T* 2.8/140 with $X=-4mm=-0.029\cdot f_{M}$ (compare zei (a)zei (b)), section 3 shows, that this assumption does not hold in general. the distance between these two parametrization planes is denoted as $F:=d-X$. To calculate the metric parametrization from a given integer parametrization, note, that the pixel pitch $\Delta_{ST}$ of the virtual sensor, i.e. the step size in the $ST$-Plane, corresponds to the microlens pitch, i.e. $\Delta_{ST}=d_{\text{ML}}$ and, as shown in figure 3, the step size in the virtual lens plane, i.e. the $UV$-plane, can be calculated by means of the triangle equality as $\Delta_{UV}=\frac{s_{px}\cdot F}{f_{m}}$ where $s_{px}$ and $f_{m}$ denote the pixel size and the microlens focal length. With these step sizes the light field parametrized in metric coordinates $(s,t,u,v)$ is given by
@@ 3 Error Analysis
$ERR_{S}(\textbf{o}):=\frac{\tilde{S}(\textbf{o})-S(\textbf{o})}{S(\textbf{o})}=X\cdot\frac{d\cdot f_{M}-\textbf{o}(d-f_{M})}{\textbf{o}\cdot f_{M}(d-X)}$ (11)
@@
$ERR_{S}(\lambda):=ERR_{S}(\lambda\cdot\textbf{o}_{f})=\frac{X(\lambda-1)}{\lambda\textbf{o}_{f}\cdot\left(\frac{X}{d}-1\right)}.$ (12)
@@
$ERR_{\tilde{\textbf{o}}}(\textbf{o}):=\frac{\tilde{\textbf{o}}(S(\textbf{o}))-\textbf{o}}{\textbf{o}}=-\frac{X(d-f_{M})+\frac{\textbf{o}}{\textbf{o}_{f}}X(f_{M}-d)+Xf_{M}d\left(\frac{1}{\textbf{o}_{f}}-\frac{1}{\textbf{o}}\right)}{f_{M}(d-X)+\frac{\textbf{o}}{\textbf{o}_{f}}X(f_{M}-d)+Xf_{M}d\frac{1}{\textbf{o}_{f}}}$ (13)
@@
$ERR_{\tilde{\textbf{o}}}(\lambda):=ERR_{\tilde{\textbf{o}}}(\lambda\textbf{o}_{f})=\frac{X(\lambda-1)^{2}}{\lambda\textbf{o}_{f}\left(1-\frac{X}{d}\right)-X(\lambda-1)\lambda}.$ (14)
@@
$ERR_{\tilde{S}}(\textbf{o}):=\frac{\textbf{o}(\tilde{S}(\textbf{o}))-\textbf{o}}{\textbf{o}}=\frac{f_{M}\mathbf{\Delta}-f_{M}X\tilde{\mathbf{\Delta}}\left(\frac{1}{\textbf{o}}-\frac{1}{\textbf{o}_{f}}\right)}{\tilde{\mathbf{\Delta}}\left(1-\frac{\textbf{o}}{\textbf{o}_{f}}\right)(f_{M}-X)+\mathbf{\Delta}f_{M}\frac{\textbf{o}}{\textbf{o}_{f}}}-1$ (15)
@@
$ERR_{\tilde{S}}(\lambda):=ERR_{\tilde{S}}(\lambda\textbf{o}_{f})=\frac{X(\lambda-1)^{2}}{\lambda\textbf{o}_{f}\left(\frac{X}{f_{M}}-1\right)-\lambda^{2}X}.$ (16)
@@
Figure 5: Left: Relative shift error based on $\lambda=\frac{\textbf{o}}{\textbf{o}_{f}}$. Mid/right: The two cases of relative object distance errors for the assumed camera focused at a finite distance which is met for a relative distance of $\frac{\textbf{o}}{\textbf{o}_{f}}=1$. Negative error values indicate an underestimation of the ground truth value while positive errors represent an overestimation.
The visualization shows that all errors diverge for $\lambda\to 0$ with a rate depending on the positional relationship between the exit pupil and the principal plane $H_{\text{cam}}$. Beyond the focus distance at $\lambda=1$ the object errors again diverge while the shift error converges according to $ERR_{S}(\lambda)\underset{\lambda\to\infty}{\longrightarrow}\frac{X}{\textbf{o}_{f}(X/d-1)}$. Note, that these graphs present an ideal refocusing case free of aliasing artifacts and limiting optical properties such as the depth of field. Hence, later experiments only verify a section of these results within the respective physical and image processing limits.
In summary these examples show a large deviation between the estimated refocusing distances in models with and without consideration of the exit pupil whenever there is a non-zero distance $X$ between the exit pupil and $H_{\text{cam}}$. This leads to the question, how prevalent a significant $X\neq 0$ is in off-the-shelf main lenses. To answer this question, the data of 866 DSLR lenses listed by Claff Claff was collected and $X$ as well as $f_{M}$ were calculated via paraxial ray-tracing for each lens.